# The holographic entanglement pattern of BTZ planar black hole from a thread perspective

### v1

@@ 4.3 Connection With Partial Entanglement Entropy Function
5 Conclusion
@@
In this paper, we explored the entanglement structure of the BTZ black hole from an intriguing perspective — that of entanglement threads. Although there have been prior studies in this direction, such as those based on bit threads or the kinematic space approach, we revisited the problem in the framework of entanglement threads, which led to subtle differences in results compared to earlier works. More importantly, this perspective offers a new interpretation of the entanglement structure of black holes, stemming from the quantum state properties associated with entanglement threads. In summary, we have pointed out that for a finite-temperature quantum state of an infinite-size CFT dual to a planar BTZ black hole, the contribution to the holographic entanglement entropy of a subregion originates not only from the complementary region within the single-sided system, but also from the other asymptotic boundary system connected through the wormhole. Within the framework of entanglement threads, we provided precise expressions for these two contributions. By incorporating the quantum state interpretation of entanglement threads, we further analyzed the apparent RT surface phase transition in the holographic BTZ black hole setup. In fact, this phase transition signifies a structure in which entanglement threads traversing the wormhole and internal entanglement within the single-sided system together form a perfect tensor pattern with quantum error-correcting properties. Our results offer a refinement of previous analyses of the partial entanglement entropy in planar BTZ black holes. While this work has focused exclusively on planar BTZ black holes, our method is, in principle, directly generalizable to the spherical BTZ black hole. Due to certain technical and interpretational subtleties that arise in this case, we leave this extension to future work.

### v2

@@ 4.3 Connection With Partial Entanglement Entropy Function
5 Conclusion and Discussion
Figure 14: Panels (a) (b) illustrate how the wormhole geometry is constructed by the quotient identification of global AdS3. Panels (a) (c) illustrate how the canonical purification of a general holographic mixed state ABAB is dual to the gluing of two copies of entanglement wedges along the minimal surfaces. Interestingly, the two perspectives may possibly be related. We leave this exploration for future work.
@@ 5 Conclusion and Discussion
In this paper, we explored the entanglement structure of the BTZ black hole from an intriguing perspective — that of entanglement threads. Although there have been prior studies in this direction, such as those based on bit threads or the kinematic space approach, we revisited the problem in the framework of entanglement threads, which led to subtle differences in results compared to earlier works. More importantly, this perspective offers a new interpretation of the entanglement structure of black holes, stemming from the quantum state properties associated with entanglement threads. In summary, we have pointed out that for a finite-temperature quantum state of an infinite-size CFT dual to a planar BTZ black hole, the contribution to the holographic entanglement entropy of a subregion originates not only from the complementary region within the single-sided system, but also from the other asymptotic boundary system connected through the wormhole. Within the framework of entanglement threads, we provided precise expressions for these two contributions. By incorporating the quantum state interpretation of entanglement threads, we further analyzed the apparent RT surface phase transition in the holographic BTZ black hole setup. In fact, this phase transition signifies a structure in which entanglement threads traversing the wormhole and internal entanglement within the single-sided system together form a perfect tensor pattern with quantum error-correcting properties. Our results offer a refinement of previous analyses of the partial entanglement entropy in planar BTZ black holes.
When the spacetime dimension of the bulk is maintained at three, our analysis on the entanglement thread structure of the planar BTZ black hole can, in principle, be generalized to the case of more general bulk spacetimes. Since a more concrete scheme requires resolving some explicit technical subtleties, we leave the detailed implementation for future work. Nevertheless, at least in principle, it is possible to articulate these ideas. This involves two key points. The first key point is that every classical solution in the AdS3 gravity (for example, various multi-boundary wormholes) could be constructed by the quotient identification of global pure AdS3 [64, 65, 66, 67, 68, 69]. As shown in Figs. 14(a) and 14(b), focusing on the static slice, for instance, the construction of the two-sided spherical BTZ geometry could be understood as the identification of the geodesics pairwise (such as the symmetric $\lambda_{1}$ and $\lambda_{2}$ in the figure) in the Poincaré disk [53, 66]. The second key point is to maintain and further develop the idea of employing tensor network states or quantum circuit representations to characterize those slices beyond the pure AdS3 case. Combining these two points, one can, according to such a recipe, use the tensor network corresponding to the original Poincaré disk to construct the holographic tensor network corresponding to other, more general slices: cut out the specified piece of the tensor network corresponding to the fundamental domain (the yellow region in the figure bounded by $\lambda_{1}$ and $\lambda_{2}$) from the original full-disk tensor network, and then glue it along those identified boundaries, i.e., contract the dangling bonds along the gluing interface. Such an idea is already familiar in previous tensor network literature, see e.g. [70, 71, 72, 73]. Then, by employing the method developed in [17], we can abstract the concept of entanglement threads within the tensor network and determine the concrete trajectories and flow fluxes of entanglement threads in the new construction. Roughly speaking, the entanglement thread flows in the original full disk are truncated at the boundary of the fundamental domain, and then will be re-glued together under the gluing operation according to some recipe, forming the new entanglement thread flows. We leave the detailed study for future work. The above line of thought can be extended to the cases of more general multi-boundary wormholes [64, 65, 66, 67, 68, 69], since we can similarly perform such an analysis by selecting the corresponding fundamental domain on the Poincaré disk and gluing along the identified surfaces.
Another interesting and direct generalization is to extend the analysis of entanglement threads for the particular finite-temperature mixed state in this paper to the analysis of more general holographic mixed states. Here we find that the ideas in [74] concerning holographic reflected entropy and its bulk dual may be extremely useful. Considering a more general bipartite holographic quantum system in an overall mixed state (such as $AB$ in Figs. 14(a) and 14(c)), [74] proposed that its canonical purification can be interpreted as dual to a CPT doubling of the original bulk geometry glued to the original geometry along minimal surfaces that separate the boundary regions. Moreover, the resulting spacetime was argued to have a continuous metric, a well-defined causal structure, and to solve Einstein’s equations with a stress-energy tensor that satisfies the Null Energy Condition. Therefore, by employing reasoning similar to that in the analysis of multi-boundary wormholes, we can cut out the specified piece of the tensor network corresponding to the bipartite mixed state, and subsequently glue it to its CPT copy along the minimal surfaces to obtain the tensor network of a new bulk spacetime. We can then analyze the entanglement thread configurations therein, which should also be formed by gluing together the local entanglement thread configurations of the two glued networks. We expect that the entanglement thread configurations in the new bulk can reasonably explain holographic reflected entropy, and be applied to analyze the intrinsic correlation structure of more general holographic mixed states. It is also an interesting direction to attempt to apply entanglement thread analysis to another concept similar to reflected entropy, namely, entanglement of purification [75, 76].
Finally, let us briefly discuss the generalization to higher dimensions. The extension of our analysis of entanglement threads for the planar BTZ black hole itself to the case of dimensions higher than three is, in principle, straightforward, because many relevant tools have already been well developed, such as the previous study of bit threads in higher dimensional spaces [10, 15] and higher dimensional kinematic space [77, 78]. Notice that our entanglement threads can be described in the language of kinematic space, where each entanglement thread can be understood as a point in kinematic space. This approach does not rely on specific dimensions, so the present analysis in principle can be directly generalized to the higher-dimensional case. However, the study of more general wormhole spacetimes and more general holographic mixed states becomes more nontrivial when extended to higher dimensions, because beyond three dimensions, generic solutions of Einstein’s equations cannot always be obtained by quotient identification of pure AdS. We leave the study of entanglement threads in various nontrivial holographic spacetimes in more general dimensions for future work.